please help prove: 1-(cos(beta)-sin(beta)^2^=sin(2beta)
Answer by Alan3354(69443)

1-(cos(beta)-sin(beta)^2^=sin(2beta)
1 - (cos^2(b) - 2sin(b)cos(b) + sin^2(b)) = 2sin(b)cos(b)
All angular arguments are the same, so assume it's b (saves typing)
----------
1 - cos^2 + 2sincos - sin^2 = 2sincos
1 - cos^2 - sin^2 = 0
sin^2 + cos^2 = 1 We all know that, right? So it's true.